In a business 30% of the shares are under government.
from the rest of the shares, 20% was distributed among the workers.
the total worth of the business is Rs.100 millions.
if the worth of the one workers share is RS.400000, how many workers are there in the company?

Answer by Theo(13342)

20% of the outstanding shares are distributed among the workers.
total worth of the business is 100 million.
one worker's shard is 400,000.

20% of 100 million is 20 million.

20 million divided by 400,000 is equal to (20 * 10^6) / (400 * 10^3).

this is equivalent to (20 * 10^6) / (4 * 10^5)

20/4 = 5
10^6 / 10^5 = 10
5 * 10 = 50

there are 50 workers in the company, as best i can figure it out.

50 workers * 400,000 per worker is equal to 50 * 4 * 10^5 which is equal to 200 * 10^5 which is equal to 20 * 10^6 which is equal to 20 million.

20 million is equal to 20% of 100 million.

solution looks good, if my assumptions of what you are looking for are correct.